Fermi National Laboratory
 

Operation Summary for April 11, 2011 - April 13, 2011

 

Monday April 11

The day shift began with Operations monitoring Tevatron (TEV) store 8635 (167.64E30) and the Recycler stash (106.2E10), with the Antiproton Source (Pbar) stacking (23.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTFB T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Linac personnel began conducting beam studies related to the Neutron Therapy Facility (NTF) at 10:22 AM. MI personnel began MI-8 collimator studies at 1:52 PM. Booster personnel tuned.

During the evening shift, Operations reported some downtime at 5:44 PM when half of the MI RF systems tripped off. Operators rebooted the low-level RF, mid-level RF, and the dampers. Operations disabled beam to T-978 at 6:36 PM. Operations terminated store 8635 at 9:49 PM. Operators prepared D0 for an access that lasted for the rest of the shift.

Tuesday April 12

The midnight shift began with operators conducting a TEV low beta magnet check, with Operations monitoring the stash (438.0E10), with Pbar stacking (20.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators conducted orbit checks. Operations put the TEV into shot setup at 2:29 AM, but that came to an abrupt halt at 3:38 AM (just before squeezing) due to an abort kicker pre-fire that caused a quench. Operations reported that cryo system recovery and kicker conditioning lasted through the end of the shift. Operations established beam to T-978 at 5:02 AM.

During the day shift, the TEV cryo system recovery continued. Operations disabled beam to T-978 at 10:13 AM; the experimenters reported that they needed to repair their detector. Operations put the TEV into shot setup at 12:11 PM and established store 8641 at 1:13 PM with an average initial luminosity (AIL) of 288.21E30. Operations reported NuMI beam permit trips due to Lambertson (I:LAM61) problems; investigations determined that the area around the power supply was too warm (over 90 degrees F). FESS personnel started working on the air conditioning (AC) system at MI-60. Operators rebooted the Recycler low-level RF system and the dampers.

During the evening shift, Operations reported that personnel inadvertently opened the wrong breaker during the MI-60 AC repairs; the MI LCW system tripped off and caused the loss of the Recycler stash. Operations recovered the MI LCW system except for the MI-20 LCW pump #2, which will need repair. Operations re-established the MI ramp and all events were learned in. At 5:18 PM, store 8641 quenched when the TEV proton abort kicker (PAK3) pre-fired again. Operations reported that investigation, recovery, and kicker conditioning lasted through to the end of the shift.

Wednesday April 13

The midnight shift began with the TEV recovering from a quench, with Operations monitoring the stash (158.0E10), with Pbar stacking (26.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators took the TEV abort kickers out of condition mode and put the TEV through a series of dry squeezes. This revealed more kicker problems; the kickers would not charge up when they were supposed to. Operators contacted TEV and Kicker personnel. TEV personnel came in to assist with the investigation. A Recycler kicker (RR20) tripped off and would not reset. Operators investigated and reported the unmistakable odor of burnt electronics. Kicker technicians came in and replaced a couple of bad resistors and an optocoupler. T-978 experimenters reported that problems with their detector continued.

The Plans for Wednesday

At 8:11 AM, Operations reported that a Linac RF station went down; Linac techs are investigating. The Run Coordinator will allow Kicker personnel to work on the TEV proton abort kicker (PAK3) and for Linac personnel to conduct NTF beam studies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/13/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory